# Scene Coordinate Reconstruction: Posing of Image Collections via Incremental Learning of a Relocalizer

### v1

@@ Appendix 0.D Baselines and Competitors
We run COLMAP in three variations: Default, fast, and “Sparse COLMAP + Reloc + BA”. Fast uses parameters recommended for large image collections beyond 1 000 images [81]. “Sparse COLMAP + Reloc + BA” uses default parameters to obtain the initial reconstruction, and fast parameters to register the remaining frames. We provide the parameters of the mapper of the default variation in Table 18. For the fast version, we provide the parameters that have been changed compared to default in Table 19. We also include the parameters of the feature extractor and the matcher shared by both variations in Tables 20 and 21.
As additional comparison to further optimize for efficiency, for the Tanks and Temples dataset we also computed in the second to last column of Tables 13, 14 and 15 a COLMAP variant “Sparse COLMAP + Reloc + BA”. As presented at the end of Section 4.3 in the main paper, this variant builds on the sparse reconstruction results, and performs the following steps:
@@
We follow the official code at https://github.com/naver/dust3r to reconstruct scenes within the 7-scene dataset from uncalibrated and unposed cameras. The DUSt3R framework cannot process all available images of a scene due to GPU memory constraints. By following the guidelines from the official repository, we configured the system to utilize a sliding window approach (w=3), enabling the accommodation of 50 frames within the 16GB memory capacity of a NVIDIA® V100 [66] GPU. As a second version, we attempted to reconstruct 50 frames using a NVIDIA® A100-40GB GPU in conjunction with the leaner 224x224 pre-trained model and an exhaustive matching configuration. We were not able to pair the full 512x512 model with exhaustive matching, even on a A100-40GB. Of the two configurations that fit into memory, the second one yields the most accurate poses under our experimental settings. Thus, we report the superior results in Table 1 of our main paper. For further insight, we have compiled the experimental statistics in Table 11.
@@
Table 11: DUSt3R on 7-Scenes S.W. denotes the sliding window configuration (window=3). E.M. denotes the exhaustive matching configuration. The performance is evaluated in PSNR (dB) using the same experiments described in Section 4.1 of the main paper. In the bottom part of the table, we show the configuration, GPU hardware, GPU memory, and reconstruction time for 50 frames on each scene.
@@
We also reconstructed scenes from the Tanks and Temples dataset using Reality Capture®, as discussed in Section 4.3 in the main paper. In Table 13, we extend the results presented in the table in Table 4 in the paper with results for training videos and intermediate videos (third column on the right side). For completeness, we also computed SSIM [99] and LPIPS [106] scores in Tables 14 and 15 respectively. On the sparse images (150-500), Reality Capture performs one order of magnitude faster than COLMAP but with slightly worse pose quality. When we ran it on Tanks and Temples with thousands of frames, Reality Capture produced a high number of disconnected components. We use the largest component as basis for our evaluation and regard all other frames as missing, as described in Appendix 0.B.
@@
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [92]   (ours) (fast) Capture [92]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 24.0 21.2 19.0 16.5 12.2k 24.4 16.9 13.5 17.7 26.3 25.2 Caterpillar 383 17.1 15.9 16.6 16.9 11.4k 18.6 17.9 18.9 18.6 18.7 18.8 Church 507 18.3 17.6 14.3 17.2 19.3k 12.1 - 11.5 16.5 18.5 18.1 Ignatius 264 20.1 17.7 17.8 19.8 7.8k 20.8 18.6 19.1 20.7 20.9 20.6 Meetingroom 371 18.6 18.1 15.6 18.0 11.1k 19.4 18.2 17.1 16.6 20.8 20.4 Truck 251 21.1 19.0 18.3 20.1 7.5k 23.6 19.1 20.6 23.0 23.4 23.3 Average 364 19.9 18.2 16.9 18.1 14.6k 19.8 18.2 16.8 18.9 21.4 21.1 Avg Time 1h 3min 5min 1.1h 74h 14h 18min 2.2h 8h 1.8h Intermediate Family 152 19.5 18.8 17.6 19.0 4.4k 21.2 19.8 19.8 18.0 21.3 21.0 Francis 302 21.6 20.7 20.7 20.1 7.8k 19.9 20.4 21.8 21.7 22.5 22.4 Horse 151 19.2 19.0 16.3 19.5 6.0k 21.6 20.7 19.2 21.7 22.6 21.4 Lighthouse 309 16.6 16.5 13.6 17.5 8.3k 19.0 16.6 18.9 18.6 19.5 19.7 Playground 307 19.1 19.2 11.4 18.7 7.7k 17.9 16.5 11.3 20.4 21.2 20.9 Train 301 16.8 15.4 13.8 16.2 12.6k 19.6 14.4 15.6 18.5 19.8 19.2 Average 254 18.8 18.3 15.6 18.5 7.8k 19.9 18.1 17.8 19.8 21.1 20.8 Avg Time 32min 2min 3min 1.3h 48h 11h 14min 2.2h 5h 1h Advanced Auditorium 302 19.6 12.2 16.7 18.7 13.6k 13.7 - 16.6 20.0 21.4 20.1 Ballroom 324 16.3 18.3 13.1 17.9 10.8k 17.2 - 10.4 18.9 18.0 17.3 Courtroom 301 18.2 17.2 12.3 17.1 12.6k 14.6 - 10.2 16.3 18.7 18.1 Palace 509 14.2 11.7 10.8 10.7 21.9k 13.8 - 8.6 11.0 15.3 12.1 Temple 302 18.1 15.7 11.8 9.7 17.5k 13.3 - 11.9 14.8 19.6 17.2 Average 348 17.3 15.0 12.9 14.8 15.6k 14.5 - 11.5 16.2 18.6 16.9 Avg Time 1h 2min 4min 1h 71h 27min 2.8h 10h 2.1h
Table 13: Tanks and Temples - PSNR ($\uparrow$). We show the pose accuracy via view synthesis with Nerfacto [91] as PSNR ($\uparrow$) in dB, and the reconstruction time. We color code results compared to COLMAP, default and fast, respectively: $>0.5$ dB better, within ±$\pm 0.5$ dB, $>0.5$ dB worse, $>1$ dB worse. Method needs sequential inputs.
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [92]   (ours) (fast) Capture [92]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 0.77 0.61 0.61 0.54 12.2k 0.77 0.56 0.49 0.58 0.84 0.79 Caterpillar 383 0.63 0.49 0.51 0.57 11.4k 0.72 0.60 0.62 0.67 0.72 0.71 Church 507 0.64 0.56 0.41 0.57 19.3k 0.29 - 0.34 0.54 0.66 0.63 Ignatius 264 0.69 0.46 0.50 0.65 7.8k 0.77 0.58 0.61 0.75 0.78 0.74 Meetingroom 371 0.67 0.62 0.48 0.63 11.1k 0.67 0.63 0.55 0.54 0.76 0.73 Truck 251 0.76 0.60 0.59 0.72 7.5k 0.83 0.59 0.69 0.80 0.83 0.82 Average 364 0.69 0.56 0.52 0.61 14.6k 0.68 0.59 0.55 0.65 0.76 0.74 Avg Time 1h 3min 5min 1.1h 74h 14h 18min 2.2h 8h 1.8h Intermediate Family 152 0.77 0.62 0.58 0.68 4.4k 0.83 0.66 0.67 0.63 0.84 0.80 Francis 302 0.84 0.74 0.77 0.74 7.8k 0.77 0.71 0.80 0.79 0.86 0.84 Horse 151 0.77 0.70 0.58 0.74 6.0k 0.79 0.72 0.69 0.82 0.85 0.79 Lighthouse 309 0.65 0.56 0.52 0.62 8.3k 0.70 0.59 0.68 0.68 0.72 0.71 Playground 307 0.54 0.51 0.23 0.50 7.7k 0.53 0.45 0.32 0.62 0.65 0.63 Train 301 0.62 0.45 0.39 0.54 12.6k 0.73 0.41 0.51 0.63 0.74 0.70 Average 254 0.70 0.60 0.51 0.64 7.8k 0.73 0.59 0.61 0.70 0.78 0.74 Avg Time 32min 2min 3min 1.3h 48h 11h 14min 2.2h 5h 1h Advanced Auditorium 302 0.74 0.39 0.58 0.66 13.6k 0.50 - 0.58 0.71 0.80 0.72 Ballroom 324 0.46 0.47 0.30 0.51 10.8k 0.48 - 0.25 0.57 0.54 0.49 Courtroom 301 0.61 0.50 0.32 0.52 12.6k 0.43 - 0.32 0.52 0.67 0.60 Palace 509 0.54 0.43 0.41 0.34 21.9k 0.50 - 0.38 0.37 0.57 0.46 Temple 302 0.68 0.52 0.43 0.22 17.5k 0.44 - 0.45 0.50 0.76 0.62 Average 348 0.61 0.46 0.41 0.45 15.6k 0.47 - 0.39 0.53 0.67 0.58 Avg Time 1h 2min 4min 1h 71h 27min 2.8h 10h 2.1h
@@
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [92]   (ours) (fast) Capture [92]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 0.21 0.36 0.42 0.51 12.2k 0.19 0.53 0.63 0.42 0.13 0.18 Caterpillar 383 0.29 0.39 0.38 0.32 11.4k 0.18 0.29 0.26 0.23 0.18 0.20 Church 507 0.31 0.38 0.64 0.38 19.3k 0.79 - 0.94 0.39 0.26 0.31 Ignatius 264 0.21 0.34 0.35 0.25 7.8k 0.15 0.27 0.25 0.16 0.14 0.17 Meetingroom 371 0.33 0.36 0.60 0.38 11.1k 0.31 0.37 0.50 0.45 0.21 0.26 Truck 251 0.18 0.31 0.34 0.22 7.5k 0.11 0.37 0.24 0.14 0.11 0.12 Average 364 0.26 0.36 0.46 0.34 14.6k 0.29 0.37 0.47 0.30 0.17 0.20 Avg Time 1h 3min 5min 1.1h 74h 14h 18min 2.2h 8h 1.8h Intermediate Family 152 0.17 0.25 0.32 0.24 4.4k 0.12 0.21 0.23 0.27 0.11 0.13 Francis 302 0.15 0.23 0.23 0.28 7.8k 0.24 0.23 0.17 0.18 0.11 0.12 Horse 151 0.21 0.23 0.41 0.23 6.0k 0.17 0.21 0.27 0.14 0.11 0.17 Lighthouse 309 0.36 0.43 0.57 0.37 8.3k 0.29 0.45 0.32 0.34 0.27 0.30 Playground 307 0.45 0.41 0.72 0.48 7.7k 0.45 0.55 0.82 0.34 0.32 0.34 Train 301 0.29 0.46 0.60 0.37 12.6k 0.20 0.64 0.45 0.29 0.18 0.21 Average 254 0.27 0.34 0.47 0.33 7.8k 0.24 0.38 0.38 0.26 0.18 0.21 Avg Time 32min 2min 3min 1.3h 48h 11h 14min 2.2h 5h 1h Advanced Auditorium 302 0.31 0.76 0.60 0.43 13.6k 0.75 - 0.64 0.37 0.21 0.33 Ballroom 324 0.43 0.38 0.68 0.39 10.8k 0.44 - 0.92 0.33 0.34 0.39 Courtroom 301 0.34 0.42 0.70 0.44 12.6k 0.61 - 0.79 0.43 0.24 0.32 Palace 509 0.56 0.77 0.87 0.78 21.9k 0.60 - 0.84 0.72 0.51 0.68 Temple 302 0.30 0.53 0.75 0.82 17.5k 0.65 - 0.70 0.53 0.21 0.39 Average 348 0.39 0.57 0.72 0.57 15.6k 0.61 - 0.78 0.47 0.30 0.42 Avg Time 1h 2min 4min 1h 71h 27min 2.8h 10h 2.1h
@@
 Pseudo GT  DROID-SLAM  BARF  NoPe-NeRF  ACE0 (COLMAP) [92] [53] [10] (ours) Bicycle 0.68 0.20 0.22 0.24 0.46 Bonsai 0.89 0.18 0.31 0.36 0.81 Counter 0.82 0.25 0.26 0.27 0.73 Garden 0.86 0.38 0.27 0.26 0.73 Kitchen 0.87 0.28 0.34 0.33 0.76 Room 0.91 0.34 0.36 0.38 0.61 Stump 0.40 0.23 0.23 0.21 0.38 Average 0.78 0.26 0.28 0.30 0.64
@@
 Pseudo GT  DROID-SLAM  BARF  NoPe-NeRF  ACE0 (COLMAP) [92] [53] [10] (ours) Bicycle 0.20 0.80 0.82 0.85 0.34 Bonsai 0.06 0.74 0.86 0.66 0.14 Counter 0.10 0.64 0.83 0.76 0.15 Garden 0.08 0.38 0.84 0.84 0.14 Kitchen 0.06 0.71 0.84 0.74 0.13 Room 0.05 0.85 0.83 0.59 0.42 Stump 0.45 0.85 0.79 0.81 0.37 Average 0.14 0.71 0.83 0.75 0.24

### v2

@@ Appendix 0.D Baselines and Competitors
We run COLMAP in three variations: Default, fast, and “Sparse COLMAP + Reloc + BA”. Fast uses parameters recommended for large image collections beyond 1 000 images [83]. “Sparse COLMAP + Reloc + BA” uses default parameters to obtain the initial reconstruction, and fast parameters to register the remaining frames. We provide the parameters of the mapper of the default variation in Table 17. For the fast version, we provide the parameters that have been changed compared to default in Table 18. We also include the parameters of the feature extractor and the matcher shared by both variations in Tables 19 and 20.
As additional comparison to further optimize for efficiency, for the Tanks and Temples dataset we also computed in the second to last column of Tables 4, 13 and 14 a COLMAP variant “Sparse COLMAP + Reloc + BA”. As presented at the end of Section 4.3, this variant builds on the sparse reconstruction results, and performs the following steps:
@@
We follow the official code at https://github.com/naver/dust3r to reconstruct scenes within the 7-Scenes dataset from uncalibrated and unposed cameras. When comparing to DUSt3R, we run ACE0 without the initialisation from a standard Kinect v1 focal length. That is, like DUSt3R, ACE0 estimates the focal length from scratch in the last column of Table 1. The DUSt3R framework cannot process all available images of a scene due to GPU memory constraints. By following the guidelines from the official repository, we configured the system to utilize a sliding window approach (w=3), enabling the accommodation of 50 frames within the 16GB memory capacity of a NVIDIA® V100 [68] GPU. As a second version, we attempted to reconstruct 50 frames using a NVIDIA® A100-40GB GPU in conjunction with the leaner 224x224 pre-trained model and an exhaustive matching configuration. We were not able to pair the full 512x512 model with exhaustive matching, even on a A100-40GB. Of the two configurations that fit into memory, the second one yields the most accurate poses under our experimental settings. Thus, we report the superior results in Table 1. For further insight, we have compiled the experimental statistics in Table 11.
@@
Table 11: DUSt3R on 7-Scenes S.W. denotes the sliding window configuration (window=3). E.M. denotes the exhaustive matching configuration. The performance is evaluated in PSNR (dB) using the same experiments described in Section 4.1. In the bottom part of the table, we show the configuration, GPU hardware, GPU memory, and reconstruction time for 50 frames on each scene.
@@
We also reconstructed scenes from the Tanks and Temples dataset using Reality Capture®, as discussed in Section 4.3. For completeness, we also computed SSIM [102] and LPIPS [109] scores in Tables 13 and 14 respectively. On the sparse images (150-500), Reality Capture performs one order of magnitude faster than COLMAP but with slightly worse pose quality. When we ran it on Tanks and Temples with thousands of frames, Reality Capture produced a high number of disconnected components. We use the largest component as basis for our evaluation and regard all other frames as missing, as described in Appendix 0.B.
@@
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [94]   (ours) (fast) Capture [94]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 0.77  0.61  0.61  0.54 12.2k 0.77  0.56  0.49  0.58  0.84  0.79 Caterpillar 383 0.63  0.49  0.51  0.57 11.4k 0.72  0.60  0.62  0.67  0.72  0.70 Church 507 0.64  0.56  0.41  0.57 19.3k 0.29  -  0.34  0.54  0.66  0.59 Ignatius 264 0.69  0.46  0.50  0.65 7.8k 0.77  0.58  0.61  0.75  0.78  0.75 Meetingroom 371 0.67  0.62  0.48  0.63 11.1k 0.67  0.63  0.55  0.54  0.76  0.72 Truck 251 0.76  0.60  0.59  0.72 7.5k 0.83  0.59  0.69  0.80  0.83  0.81 Average 364 0.69  0.56  0.52  0.61 14.6k 0.68  0.59  0.55  0.65  0.76  0.73 Avg Time 1h  3min  5min  1.1h 74h  14h  18min  2.2h  8h  1.8h Intermediate Family 152 0.77  0.62  0.58  0.68 4.4k 0.83  0.66  0.67  0.63  0.84  0.82 Francis 302 0.84  0.74  0.77  0.74 7.8k 0.77  0.71  0.80  0.79  0.86  0.83 Horse 151 0.77  0.70  0.58  0.74 6.0k 0.79  0.72  0.69  0.82  0.85  0.83 Lighthouse 309 0.65  0.56  0.52  0.62 8.3k 0.70  0.59  0.68  0.68  0.72  0.73 Playground 307 0.54  0.51  0.23  0.50 7.7k 0.53  0.45  0.32  0.62  0.65  0.63 Train 301 0.62  0.45  0.39  0.54 12.6k 0.73  0.41  0.51  0.63  0.74  0.63 Average 254 0.70  0.60  0.51  0.64 7.8k 0.73  0.59  0.61  0.70  0.78  0.75 Avg Time 32min  2min  3min  1.3h 48h  11h  14min  2.2h  5h  1h Advanced Auditorium 302 0.74  0.39  0.58  0.66 13.6k 0.50  -  0.58  0.71  0.80  0.70 Ballroom 324 0.46  0.47  0.30  0.51 10.8k 0.48  -  0.25  0.57  0.54  0.41 Courtroom 301 0.61  0.50  0.32  0.52 12.6k 0.43  -  0.32  0.52  0.67  0.59 Palace 509 0.54  0.43  0.41  0.34 21.9k 0.50  -  0.38  0.37  0.57  0.46 Temple 302 0.68  0.52  0.43  0.22 17.5k 0.44  -  0.45  0.50  0.76  0.58 Average 348 0.61  0.46  0.41  0.45 15.6k 0.47  -  0.39  0.53  0.67  0.55 Avg Time 1h  2min  4min  1h 71h  27min  2.8h  10h  2.1h
@@
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [94]   (ours) (fast) Capture [94]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 0.21  0.36  0.42  0.51 12.2k 0.19  0.53  0.63  0.42  0.13  0.18 Caterpillar 383 0.29  0.39  0.38  0.32 11.4k 0.18  0.29  0.26  0.23  0.18  0.19 Church 507 0.31  0.38  0.64  0.38 19.3k 0.79  -  0.94  0.39  0.26  0.36 Ignatius 264 0.21  0.34  0.35  0.25 7.8k 0.15  0.27  0.25  0.16  0.14  0.16 Meetingroom 371 0.33  0.36  0.60  0.38 11.1k 0.31  0.37  0.50  0.45  0.21  0.27 Truck 251 0.18  0.31  0.34  0.22 7.5k 0.11  0.37  0.24  0.14  0.11  0.13 Average 364 0.26  0.36  0.46  0.34 14.6k 0.29  0.37  0.47  0.30  0.17  0.22 Avg Time 1h  3min  5min  1.1h 74h  14h  18min  2.2h  8h  1.8h Intermediate Family 152 0.17  0.25  0.32  0.24 4.4k 0.12  0.21  0.23  0.27  0.11  0.13 Francis 302 0.15  0.23  0.23  0.28 7.8k 0.24  0.23  0.17  0.18  0.11  0.13 Horse 151 0.21  0.23  0.41  0.23 6.0k 0.17  0.21  0.27  0.14  0.11  0.13 Lighthouse 309 0.36  0.43  0.57  0.37 8.3k 0.29  0.45  0.32  0.34  0.27  0.26 Playground 307 0.45  0.41  0.72  0.48 7.7k 0.45  0.55  0.82  0.34  0.32  0.33 Train 301 0.29  0.46  0.60  0.37 12.6k 0.20  0.64  0.45  0.29  0.18  0.30 Average 254 0.27  0.34  0.47  0.33 7.8k 0.24  0.38  0.38  0.26  0.18  0.21 Avg Time 32min  2min  3min  1.3h 48h  11h  14min  2.2h  5h  1h Advanced Auditorium 302 0.31  0.76  0.60  0.43 13.6k 0.75  -  0.64  0.37  0.21  0.37 Ballroom 324 0.43  0.38  0.68  0.39 10.8k 0.44  -  0.92  0.33  0.34  0.46 Courtroom 301 0.34  0.42  0.70  0.44 12.6k 0.61  -  0.79  0.43  0.24  0.34 Palace 509 0.56  0.77  0.87  0.78 21.9k 0.60  -  0.84  0.72  0.51  0.67 Temple 302 0.30  0.53  0.75  0.82 17.5k 0.65  -  0.70  0.53  0.21  0.44 Average 348 0.39  0.57  0.72  0.57 15.6k 0.61  -  0.78  0.47  0.30  0.46 Avg Time 1h  2min  4min  1h 71h  27min  2.8h  10h  2.1h
@@
 Pseudo GT  DROID-SLAM  BARF  NoPe-NeRF  ACE0 (COLMAP) [94] [55] [10] (ours) Bicycle 0.68 0.20 0.22 0.24 0.47 Bonsai 0.89 0.18 0.31 0.36 0.83 Counter 0.82 0.25 0.26 0.27 0.77 Garden 0.86 0.38 0.27 0.26 0.77 Kitchen 0.87 0.28 0.34 0.33 0.80 Room 0.91 0.34 0.36 0.38 0.61 Stump 0.40 0.23 0.23 0.21 0.40 Average 0.78 0.26 0.28 0.30 0.67
@@
 Pseudo GT  DROID-SLAM  BARF  NoPe-NeRF  ACE0 (COLMAP) [94] [55] [10] (ours) Bicycle 0.20 0.80 0.82 0.85 0.33 Bonsai 0.06 0.74 0.86 0.66 0.11 Counter 0.10 0.64 0.83 0.76 0.12 Garden 0.08 0.38 0.84 0.84 0.12 Kitchen 0.06 0.71 0.84 0.74 0.11 Room 0.05 0.85 0.83 0.59 0.43 Stump 0.45 0.85 0.79 0.81 0.36 Average 0.14 0.71 0.83 0.75 0.23